# A Two-Timescale Primal-Dual Framework for Reinforcement Learning via Online Dual Variable Guidance

### v2

@@ 4.2 Convergence Rate Analysis of Algorithm 2
The rate of convergence of the order $\widetilde{\mathcal{O}}(k^{-2/3})$ recovers the existing rate of Zeng et al. [2024] for two-timescale stochastic gradient schemes under Markovian model access, albeit under the more challenging setting of buffer-based biased gradient estimates and accounting for asynchronous updating with visitation-adjusted stepsizes.
Using Lipschitz continuity of $\lambda(\cdot)$ and $\rho\mapsto\pi_{\rho}$ on the compact set $H$, the result further implies $\widetilde{\mathcal{O}}(k^{-2/3})$ last-iterate mean-square convergence of the primal best response $\lambda(\rho_{k})$ to the optimal regularized value $V_{r}^{\star}=\lambda(\rho^{\star})$, and of the induced policy $\pi_{\rho_{k}}$ to $\pi_{r}^{\star}$.
The proof of Theorem 4.2 now builds on two contraction inequalities, one each for the primal and the dual variable, which are then connected by the two-timescale lemma [Zeng et al., 2024, Lemma 4]. Recall that we overload the Lipschitz constant notation and denote all Lipschitz constants by the envelope $L$.
Proposition 4.3.
For $\delta\in(0,1)$ let $\mathcal{G}_{\delta}$ denote the high-probability visitation event of Proposition 4.2. Let $C_{1}$ be as in (25). Then, for all $k\geq\max\{K(\delta),\mathcal{K}\}$, the primal variable tracking error $\mathbb{E}\left[\|V_{k}-\lambda(\rho_{k})\|_{2}^{2}|\mathcal{G}_{\delta}\right]$ satisfies
@@
Proposition 4.4.
For $\delta\in(0,1)$ let $\mathcal{G}_{\delta}$ denote the high-probability visitation event of Proposition 4.2. Then for all $k\geq\max\{K(\delta),\,\mathcal{K}\}$, the reduced objective error $\mathbb{E}[f(\rho^{\star})-f(\rho_{k})|\mathcal{G}_{\delta}]$ satisfies
$\mathbb{E}[f(\rho^{\star})-f(\rho_{k+1})|\mathcal{G}_{\delta}]$ $\leq(1-\mu_{\mathrm{opt}}\,\beta_{k})\mathbb{E}[f(\rho^{\star})-f(\rho_{k})|\mathcal{G}_{\delta}]+\frac{L^{2}\beta_{\mathrm{env},k}}{2}\mathbb{E}[\|V_{k}-\lambda(\rho_{k})\|_{2}^{2}|\mathcal{G}_{\delta}]$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +\frac{25L^{2}B^{3}}{2}\tau_{k}^{2}\beta_{\mathrm{env},k-\tau_{k}}\beta_{\mathrm{env},k},$
@@
The proof of Theorem 4.2 now consists of combining the recursive contractions from Propositions 4.3 and 4.4 using the two-timescale lemma [Zeng et al., 2024, Lemma 4], restated here for completeness.
Lemma 4.2.
(Two-timescale Lemma, [Zeng et al., 2024, Lemma 4, Case 1])
Let $\left\{a_{k},b_{k},c_{k},d_{k},e_{k},f_{k}\right\}$ be non-negative sequences satisfying $\frac{a_{k+1}}{d_{k+1}}\leq\frac{a_{k}}{d_{k}}<1$ for all $k\geq 0$. Let $\left\{x_{k}\right\},\left\{y_{k}\right\}$ be two nonnegative sequences. Suppose that $x_{k},y_{k}$ satisfy the following coupled inequalities:
@@
Proof of Theorem 4.2
Fix $\delta\in(0,1)$ and work on $\mathcal{G}_{\delta}$ defined in Proposition 4.2. Let $k_{0}:=\max\{K(\delta),\mathcal{K}\}$ be the burn-in index such that the visitation floor bounds and the technical conditions in (24) hold for all $k\geq k_{0}$. Set
@@
(I.) Following the two coupled inequalities of Proposition 4.3 and Proposition 4.4 it holds that for all $k\geq k_{0}$
@@
$a_{k}:=\mu_{\mathrm{opt}}\beta_{k},\qquad b_{k}:=\frac{L^{2}}{2}\beta_{\mathrm{env},k},\qquad c_{k}:=\frac{25L^{2}B^{3}}{2}\,\tau_{k}^{2}\,\beta_{\mathrm{env},k-\tau_{k}}\beta_{\mathrm{env},k},$
@@
Inserting the stepsize sequences into $r_{k}(\delta)$ (see Proposition 4.3) and adding the constant factors, we find that, with constant $C_{R}(\delta):=\tfrac{12}{p_{\star}\tilde{\eta}_{V}}C_{\mathrm{buf}}(\delta)^{2}|\mathcal{S}|\alpha_{0}+\tfrac{12}{p_{\star}^{2}}\alpha_{0}^{2}(\tfrac{6C_{1}C_{\tau}^{2}}{c_{\tau}}+2D^{2})+\tfrac{16}{p_{\star}^{2}}L^{2}B^{2}\beta_{0}^{2}(1+\tfrac{2}{\alpha_{0}\tilde{\eta}_{V}})$, for all $k\geq k_{0}$,
@@
(II.) We proceed to determine a feasible constant $A\in\mathbb{R}$ to apply Lemma 4.2. Let $k\geq k_{0}$. Inserting the chosen stepsize scheme (21) into the fraction $a_{k}/d_{k}$ gives
$\frac{a_{k}}{d_{k}}=\frac{\mu_{\mathrm{opt}}\beta_{k}}{(\tilde{\eta}_{V}/4)\alpha_{k}}=\frac{4\mu_{\mathrm{opt}}}{\tilde{\eta}_{V}}\frac{\beta_{k}}{\alpha_{k}}=\frac{4\mu_{\mathrm{opt}}}{\tilde{\eta}_{V}}\frac{\beta_{0}}{\alpha_{0}(k+1)^{1/3}}.$ (29)
By the imposed conditions on the stepsize parameters, $\beta_{0}\geq\tfrac{2}{\mu_{\mathrm{opt}}}$ and $\alpha_{0}$ is set such that $\frac{\beta_{0}}{\alpha_{0}}\leq\frac{\tilde{\eta}_{V}}{8\mu_{\mathrm{opt}}}$. Since $\tfrac{\beta_{k}}{\alpha_{k}}=\tfrac{\beta_{0}}{\alpha_{0}(k+1)^{1/3}}$ is non-increasing in $k$, this implies $\tfrac{a_{k}}{d_{k}}\leq\tfrac{1}{2}(k+1)^{-1/3}\leq\tfrac{1}{2}$. Moreover, by the stepsize envelope construction (see (23)),
@@
We set $A:=\frac{4L^{2}}{p_{\star}\mu_{\mathrm{opt}}}.$ Then, using $\tfrac{a_{k}}{d_{k}}\leq\tfrac{1}{2}$ and $b_{k}=\tfrac{L^{2}}{2}\beta_{\mathrm{env},k}\leq\tfrac{2L^{2}}{p_{\star}}\beta_{k}$, we obtain
$Aa_{k}\Bigl(1-\frac{a_{k}}{d_{k}}\Bigr)-b_{k}\geq\frac{A}{2}a_{k}-b_{k}\geq\frac{2L^{2}}{p_{\star}\mu_{\mathrm{opt}}}\mu_{\mathrm{opt}}\beta_{k}-\frac{2L^{2}}{p_{\star}}\beta_{k}=0,$
and $Ae_{k}/d_{k}=0$ since $e_{k}=0$. Hence, Lemma 4.2 applies for the coupled system (26)–(27) and gives
@@
$A\frac{a_{k}}{d_{k}}f_{k}\leq\frac{2L^{2}C_{R}(\delta)}{p_{\star}\mu_{\mathrm{opt}}}\cdot\frac{\log^{2}(k+1)}{(k+1)^{5/3}}.$
@@
$c_{k}+A\frac{a_{k}}{d_{k}}f_{k}\leq C_{2}(\delta)\frac{\log^{2}(k+1)}{(k+1)^{5/3}},\qquad C_{2}(\delta):=\frac{2L^{2}C_{R}(\delta)}{p_{\star}\mu_{\mathrm{opt}}}+\frac{200L^{2}B^{3}C_{\tau}^{2}\beta_{0}^{2}}{c_{\tau}p_{\star}^{2}},$ (31)
@@
$\prod_{t=\ell+1}^{k-1}\Bigl(1-\frac{a_{t}}{2}\Bigr)\leq\exp\Bigl(-\frac{1}{2}\sum_{t=\ell+1}^{k-1}a_{t}\Bigr)=\exp\Bigl(-\frac{\mu_{\mathrm{opt}}\beta_{0}}{2}\sum_{t=\ell+1}^{k-1}\frac{1}{t+1}\Bigr)\leq\Bigl(\frac{\ell+2}{k+1}\Bigr)^{\tfrac{\mu_{\mathrm{opt}}\beta_{0}}{2}}\leq\frac{\ell+2}{k+1},$
where the last inequality is due to the condition $\beta_{0}\geq\tfrac{2}{\mu_{\mathrm{opt}}}$. With the same reasoning $\prod_{t={k_{0}}}^{k-1}\left(1-\frac{a_{t}}{2}\right)\leq\tfrac{k_{0}+1}{k+1}$. Inserting the harmonic sum bounds and (31) in (30) and using $\frac{\ell+2}{k+1}\leq 2\tfrac{\ell+1}{k+1}$ yields, for all $k\geq k_{0}$,
$x_{k}\leq\left(x_{k_{0}}+\frac{16L^{2}}{p_{\star}\tilde{\eta}_{V}}\frac{\beta_{k_{0}}}{\alpha_{k_{0}}}y_{k_{0}}\right)\frac{k_{0}+1}{k+1}+\frac{2C_{2}(\delta)\log^{2}(k+1)}{k+1}\sum_{\ell=k_{0}}^{k-1}\frac{1}{(\ell+1)^{2/3}},$
@@
$\mathbb{E}\left[f(\rho^{\star})-f(\rho_{k})|\mathcal{G}_{\delta}\right]\leq\big(\mathbb{E}\left[f(\rho^{\star})-f(\rho_{k_{0}})|\mathcal{G}_{\delta}\right]$ $+\frac{16L^{2}}{p_{\star}\tilde{\eta}_{V}}\frac{\beta_{k_{0}}}{\alpha_{k_{0}}}\mathbb{E}\left[\|V_{k_{0}}-\lambda(\rho_{k_{0}})\|_{2}^{2}|\mathcal{G}_{\delta}\right]\big)\frac{k_{0}+1}{k+1}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ +\frac{6C_{2}(\delta)\log^{2}(k+1)}{(k+1)^{2/3}}.$
Since $\rho^{\star}\in\mathrm{int}H$, it holds that $\nabla f(\rho^{\star})=0$, hence with the strong concavity of $f$ (see Proposition 4.2-(iv) and the proof in Lemma C.5)
@@ 5 Numerical Results
We evaluate PGDA-RL on the FrozenLake-v1 environment [Towers et al., 2024], a standard reinforcement learning benchmark with discrete states and actions. [^4] All experiments are implemented in Python, and the code for reproducing all numerical results is available from https://github.com/AxelFW/tt-primal-dual-rl. The environment consists of a $4\times 4$ grid in which the agent must reach a goal state while avoiding holes. The action space contains four actions (up, down, left, right) and transitions are stochastic due to the “slippery” dynamics. To adapt this episodic task to our infinite-horizon setting, we treat reaching the goal or a hole as a transition back to the start state, allowing continuous learning. The reward for reaching the goal is $100$, and all other transitions yield zero reward.
We set the model parameters to $\{\gamma,\eta_{\rho},\eta_{V}\}=\{0.9,0.1,0.1\}$. The asynchronous stepsizes are $\alpha_{k}=(1+\tilde{k})^{-2/3}$ and $\beta_{k}=(1+\tilde{k})^{-1}$ with $\tilde{k}=9+k/100$. We apply Algorithm 2 with on-policy exploration and an additional $\epsilon$-greedy strategy with linearly decaying $\epsilon$ ($\epsilon_{0},\epsilon_{K}]=[1,0.1]$, $K=10^{5}$) to mitigate infrequent updates of rarely visited statereplay-buffer biasaction pairs. Replay buffers are capped at size $10^{3}$. In the numerical implementation, we additionally impose a lower bound of $10^{-12}$ on the entries of the dual iterate after projection to avoid numerical instabilities when projected components become extremely small.
@@
Figure 1: Value function and policy evaluation during training. The plots show the mean (solid lines) and shaded $\pm 2$ standard errors over $10$ independent training runs. (a) rRMSE to the optimal regularized value. (b) Initial-state value under the dual-iterate policy. (c) rRMSE to the optimal (unregularized) value and comparison to the generator-based benchmark SPD-dMDP [Chen and Wang, 2016]. (d) KL divergence between the learned policy $\pi_{\rho_{k}}$ and the optimal regularized policy $\pi_{r}^{\star}$.
Figure 1 summarizes the learning dynamics. Panel (a) shows that both the primal value iterate and the value induced by the current dual policy steadily reduce their relative value error on non-terminal states. The dual-policy value improves faster and reaches a lower error level over the shown horizon. This separation is consistent with the two-timescale design: while $\rho_{k}$ evolves slowly toward $\rho^{\star}$, the value update primarily tracks the moving best response $\lambda(\rho_{k})$, so the policy-induced value can stabilize earlier than the primal iterate over the finite training horizon. Panel (b) tracks the initial-state value under the dual-iterate policy: after a rapid increase during early training, the curve approaches the initial state value of the optimal regularized policy. The separation between the optimal unregularized value and the value of the optimal regularized policy shows the suboptimality induced by the entropy regularization. Panel (c) compares PGDA-RL against the generator-based benchmark SPD-dMDP [Chen and Wang, 2016] that builds on the LP approach to (unregularized) MDPs. SPD-dMDP improves more quickly early on, whereas PGDA-RL closes the gap over longer training and achieves comparable final accuracy over the shown horizon. A key difference is the sampling model: SPD-dMDP draws transitions from uniformly sampled state-action pairs, while our method relies on trajectory data generated by the current behavior policy. Moreover, SPD-dMDP reports the averaged iterate as output, which yields a visibly smoother curve than the last-iterate behavior shown for PGDA-RL. Finally, panel (d) reports the KL divergence between the learned dual iterate policy and the optimal regularized policy summed uniformly over non-terminal states, $\mathsf{D}(\pi_{r}^{\star}\parallel\pi_{\rho_{k}})=\sum_{s\in\mathcal{S},a\in\mathcal{A}}\pi_{r}^{\star}(a|s)\log\left(\frac{\pi_{r}^{\star}(a|s)}{\pi_{\rho_{k}}(a|s)}\right)$, which decreases consistently over training, indicating convergence in policy space alongside value improvement. In summary, the simulations support our theoretical findings by demonstrating convergence of the dual-induced policy $\pi_{\rho_{k}}$ toward the unique optimal regularized policy $\pi_{r}^{\star}$ on all non-terminal states.

### v3

@@ 4.2 Convergence Rate Analysis of Algorithm 2
The rate of convergence of the order $\widetilde{\mathcal{O}}(k^{-2/3})$ recovers the existing rate of Zeng et al. (2024) for two-timescale stochastic gradient schemes under Markovian model access, albeit under the more challenging setting of buffer-based biased gradient estimates and accounting for asynchronous updating with visitation-adjusted stepsizes.
As a consequence of the finite-time convergence rate of the dual iterates (Theorem 16), we obtain the following bound on the unregularized value suboptimality of the dual-induced policies; see Appendix C.3 for the proof.
Corollary 17 (Value suboptimality of the induced policy)
Under the assumptions of Theorem 16, let $L_{\pi}>0$ satisfy
$\max_{s\in\mathcal{S}}\|\pi_{\rho}(\cdot|s)-\pi_{\hat{\rho}}(\cdot|s)\|_{1}\leq L_{\pi}\|\rho-\hat{\rho}\|_{2},\qquad\forall\,\rho,\hat{\rho}\in H,$
and define
$L_{V}:=\frac{C_{r}}{(1-\gamma)^{2}}L_{\pi},\qquad B_{\mathrm{reg}}:=\frac{\eta_{\rho}\log|\mathcal{A}|}{1-\gamma}.$
Then, for every $s\in\mathcal{S}$ and all $k\geq k_{0}$,
$\mathbb{E}\!\left[\left(V^{\star}_{ur}(s)-V^{\pi_{\rho_{k}}}_{ur}(s)-B_{\mathrm{reg}}\right)_{+}^{2}\;\middle|\;\mathcal{G}_{\delta}\right]\leq L_{V}^{2}\,\mathbb{E}\!\left[\|\rho_{k}-\rho^{\star}\|_{2}^{2}\;\middle|\;\mathcal{G}_{\delta}\right].$
In particular, the excess unregularized value suboptimality beyond the deterministic regularization bias $B_{\mathrm{reg}}$ inherits the conditional mean-square rate $\widetilde{\mathcal{O}}(k^{-2/3})$ from Theorem 16.
The proof of Theorem 16 builds on two contraction inequalities, one each for the primal and the dual variable, which are then connected by the two-timescale lemma (Zeng et al., 2024, Lemma 4). Recall that we overload the Lipschitz constant notation and denote all Lipschitz constants by the envelope $L$.
Proposition 18
For $\delta\in(0,1)$ let $\mathcal{G}_{\delta}$ denote the high-probability visitation event of Proposition 15. Let $C_{1}$ be as in (25). Then, for all $k\geq\max\{K(\delta),\mathcal{K}\}$, the primal variable tracking error $\mathbb{E}\left[\|V_{k}-\lambda(\rho_{k})\|_{2}^{2}|\mathcal{G}_{\delta}\right]$ satisfies
@@
Proposition 19
For $\delta\in(0,1)$ let $\mathcal{G}_{\delta}$ denote the high-probability visitation event of Proposition 15. Then for all $k\geq\max\{K(\delta),\,\mathcal{K}\}$, the reduced objective error $\mathbb{E}[f(\rho^{\star})-f(\rho_{k})|\mathcal{G}_{\delta}]$ satisfies
$\mathbb{E}[f(\rho^{\star})-f(\rho_{k+1})|\mathcal{G}_{\delta}]$ $\leq(1-p_{\star}\mu_{\mathrm{opt}}\,\beta_{k})\mathbb{E}[f(\rho^{\star})-f(\rho_{k})|\mathcal{G}_{\delta}]+\frac{L^{2}\beta_{\mathrm{env},k}}{2p_{\star}}\mathbb{E}[\|V_{k}-\lambda(\rho_{k})\|_{2}^{2}|\mathcal{G}_{\delta}]$
$~~~~~~~~~~+\frac{25L^{2}B^{3}}{2}\tau_{k}^{2}\beta_{\mathrm{env},k-\tau_{k}}\beta_{\mathrm{env},k},$
@@
The proof of Theorem 16 now consists of combining the recursive contractions from Propositions 18 and 19 using the two-timescale lemma (Zeng et al., 2024, Lemma 4, Case 1), restated here for completeness.
Lemma 20 (Two-timescale lemma, Zeng et al. 2024)
Let $\left\{a_{k},b_{k},c_{k},d_{k},e_{k},f_{k}\right\}$ be
non-negative sequences satisfying $\frac{a_{k+1}}{d_{k+1}}\leq\frac{a_{k}}{d_{k}}<1$ for all $k\geq 0$. Let $\left\{x_{k}\right\},\left\{y_{k}\right\}$ be two nonnegative sequences. Suppose that $x_{k},y_{k}$ satisfy the following coupled inequalities:
@@
Proof of Theorem 16
Fix $\delta\in(0,1)$ and work on $\mathcal{G}_{\delta}$ defined in Proposition 15. Let $k_{0}:=\max\{K(\delta),\mathcal{K}\}$ be the burn-in index such that the visitation floor bounds and the technical conditions in (24) hold for all $k\geq k_{0}$. Set
@@
(I.) Following the two coupled inequalities of Proposition 18 and Proposition 19 it holds that for all $k\geq k_{0}$
@@
$a_{k}:=p_{\star}\mu_{\mathrm{opt}}\beta_{k},\qquad b_{k}:=\frac{L^{2}}{2p_{\star}}\beta_{\mathrm{env},k},\qquad c_{k}:=\frac{25L^{2}B^{3}}{2}\,\tau_{k}^{2}\,\beta_{\mathrm{env},k-\tau_{k}}\beta_{\mathrm{env},k},$
@@
Inserting the stepsize sequences into $r_{k}(\delta)$ (see Proposition 18) and adding the constant factors, we find that, with constant $C_{R}(\delta):=\tfrac{12}{p_{\star}\tilde{\eta}_{V}}C_{\mathrm{buf}}(\delta)^{2}|\mathcal{S}|\alpha_{0}+\tfrac{12}{p_{\star}^{2}}\alpha_{0}^{2}(\tfrac{6C_{1}C_{\tau}^{2}}{c_{\tau}}+2D^{2})+\tfrac{16}{p_{\star}^{2}}L^{2}B^{2}\beta_{0}^{2}(1+\tfrac{2}{\alpha_{0}\tilde{\eta}_{V}})$, for all $k\geq k_{0}$,
@@
(II.) We proceed to determine a feasible constant $A\in\mathbb{R}$ to apply Lemma 20. Let $k\geq k_{0}$. Inserting the chosen stepsize scheme (21) into the fraction $a_{k}/d_{k}$ gives
$\frac{a_{k}}{d_{k}}=\frac{p_{\star}\mu_{\mathrm{opt}}\beta_{k}}{(\tilde{\eta}_{V}/4)\alpha_{k}}=\frac{4p_{\star}\mu_{\mathrm{opt}}}{\tilde{\eta}_{V}}\frac{\beta_{k}}{\alpha_{k}}=\frac{4p_{\star}\mu_{\mathrm{opt}}}{\tilde{\eta}_{V}}\frac{\beta_{0}}{\alpha_{0}(k+1)^{1/3}}.$ (29)
By the imposed conditions on the stepsize parameters, $\beta_{0}\geq\tfrac{2}{p_{\star}\mu_{\mathrm{opt}}}$ and $\alpha_{0}$ is set such that $\frac{\beta_{0}}{\alpha_{0}}\leq\frac{\tilde{\eta}_{V}}{8p_{\star}\mu_{\mathrm{opt}}}$. Since $\tfrac{\beta_{k}}{\alpha_{k}}=\tfrac{\beta_{0}}{\alpha_{0}(k+1)^{1/3}}$ is non-increasing in $k$, this implies $\tfrac{a_{k}}{d_{k}}\leq\tfrac{1}{2}(k+1)^{-1/3}\leq\tfrac{1}{2}$. Moreover, by the stepsize envelope construction (see 23),
@@
We set $A:=\frac{4L^{2}}{p_{\star}^{3}\mu_{\mathrm{opt}}}.$ Then, using $\tfrac{a_{k}}{d_{k}}\leq\tfrac{1}{2}$ and $b_{k}=\tfrac{L^{2}}{2p_{\star}}\beta_{\mathrm{env},k}\leq\tfrac{2L^{2}}{p_{\star}^{2}}\beta_{k}$, we obtain
$Aa_{k}\Bigl(1-\frac{a_{k}}{d_{k}}\Bigr)-b_{k}\geq\frac{A}{2}a_{k}-b_{k}\geq\frac{2L^{2}}{p_{\star}^{2}\mu_{\mathrm{opt}}}\mu_{\mathrm{opt}}\beta_{k}-\frac{2L^{2}}{p_{\star}^{2}}\beta_{k}=0,$
and $Ae_{k}/d_{k}=0$ since $e_{k}=0$. Hence, Lemma 20 applies for the coupled system (26)–(27) and gives
@@
$A\frac{a_{k}}{d_{k}}f_{k}\leq\frac{2L^{2}C_{R}(\delta)}{p_{\star}^{3}\mu_{\mathrm{opt}}}\cdot\frac{\log^{2}(k+1)}{(k+1)^{5/3}}.$
@@
$c_{k}+A\frac{a_{k}}{d_{k}}f_{k}\leq C_{2}(\delta)\frac{\log^{2}(k+1)}{(k+1)^{5/3}},\qquad C_{2}(\delta):=\frac{2L^{2}C_{R}(\delta)}{p_{\star}^{3}\mu_{\mathrm{opt}}}+\frac{200L^{2}B^{3}C_{\tau}^{2}\beta_{0}^{2}}{c_{\tau}p_{\star}^{2}},$ (31)
@@
$\prod_{t=\ell+1}^{k-1}\Bigl(1-\frac{a_{t}}{2}\Bigr)$ $\leq\exp\Bigl(-\frac{1}{2}\sum_{t=\ell+1}^{k-1}a_{t}\Bigr)=\exp\Bigl(-\frac{p_{\star}\mu_{\mathrm{opt}}\beta_{0}}{2}\sum_{t=\ell+1}^{k-1}\frac{1}{t+1}\Bigr)$
$\leq\Bigl(\frac{\ell+2}{k+1}\Bigr)^{\tfrac{p_{\star}\mu_{\mathrm{opt}}\beta_{0}}{2}}\leq\frac{\ell+2}{k+1},$
where the last inequality is due to the condition $\beta_{0}\geq\tfrac{2}{p_{\star}\mu_{\mathrm{opt}}}$. With the same reasoning $\prod_{t={k_{0}}}^{k-1}\left(1-\frac{a_{t}}{2}\right)\leq\tfrac{k_{0}+1}{k+1}$. Inserting the harmonic sum bounds and (31) in (30) and using $\frac{\ell+2}{k+1}\leq 2\tfrac{\ell+1}{k+1}$ yields, for all $k\geq k_{0}$,
$x_{k}\leq\left(x_{k_{0}}+\frac{16L^{2}}{p_{\star}^{3}\tilde{\eta}_{V}}\frac{\beta_{k_{0}}}{\alpha_{k_{0}}}y_{k_{0}}\right)\frac{k_{0}+1}{k+1}+\frac{2C_{2}(\delta)\log^{2}(k+1)}{k+1}\sum_{\ell=k_{0}}^{k-1}\frac{1}{(\ell+1)^{2/3}},$
@@
$\mathbb{E}\left[f(\rho^{\star})-f(\rho_{k})|\mathcal{G}_{\delta}\right]\leq\big(\mathbb{E}\left[f(\rho^{\star})-f(\rho_{k_{0}})|\mathcal{G}_{\delta}\right]$ $+\frac{16L^{2}}{p_{\star}^{3}\tilde{\eta}_{V}}\frac{\beta_{k_{0}}}{\alpha_{k_{0}}}\mathbb{E}\left[\|V_{k_{0}}-\lambda(\rho_{k_{0}})\|_{2}^{2}|\mathcal{G}_{\delta}\right]\big)\frac{k_{0}+1}{k+1}$
$~~~+\frac{6C_{2}(\delta)\log^{2}(k+1)}{(k+1)^{2/3}}.$
Since $\rho^{\star}\in\mathrm{int}H$, it holds that $\nabla f(\rho^{\star})=0$, hence with the strong concavity of $f$ (see Proposition 15-(iv) and the proof in Lemma 29)
@@ 5 Numerical Results
We evaluate PGDA-RL on the FrozenLake-v1 environment (Towers et al., 2024), a standard reinforcement learning benchmark with discrete states and actions. [^4] All experiments are implemented in Python, and the code for reproducing all numerical results is available from https://github.com/AxelFW/tt-primal-dual-rl. The environment consists of a $4\times 4$ grid in which the agent must reach a goal state while avoiding holes. The action space contains four actions (up, down, left, right) and transitions are stochastic due to the “slippery” dynamics. To adapt this episodic task to our infinite-horizon setting, we treat reaching the goal or a hole as a transition back to the start state, allowing continuous learning. The reward for reaching the goal is $100$, and all other transitions yield zero reward.
We set the model parameters to $\{\gamma,\eta_{\rho},\eta_{V}\}=\{0.9,0.1,0.1\}$. The asynchronous stepsizes are $\alpha_{k}=(1+\tilde{k})^{-2/3}$ and $\beta_{k}=(1+\tilde{k})^{-1}$ with $\tilde{k}=9+k/100$. We apply Algorithm 2 with on-policy exploration and an additional $\epsilon$-greedy strategy with linearly decaying $\epsilon$ ($\epsilon_{0},\epsilon_{K}]=[1,0.1]$, $K=10^{5}$) to mitigate infrequent updates of rarely visited state-action pairs. Replay buffers are capped at size $10^{3}$. In the numerical implementation, we additionally impose a lower bound of $10^{-12}$ on the entries of the dual iterate after projection to avoid numerical instabilities when projected components become extremely small.
@@
Figure 1: Value function and policy evaluation during training. The plots show the mean (solid lines) and shaded $\pm 2$ standard errors over 1010 independent training runs. (a) rRMSE to the optimal regularized value. (b) Initial-state value under the dual-iterate policy. (c) rRMSE to the optimal (unregularized) value and comparison to the generator-based benchmark SPD-dMDP (Chen and Wang, 2016). (d) KL-divergence between the learned policy $\pi_{\rho_{k}}$ and the optimal regularized policy $\pi_{r}^{\star}$.
Figure 1 summarizes the learning dynamics. Panel (a) shows that both the primal value iterate and the value induced by the current dual policy steadily reduce their relative value error on non-terminal states. The dual-policy value improves faster and reaches a lower error level over the shown horizon. This separation is consistent with the two-timescale design: while $\rho_{k}$ evolves slowly toward $\rho^{\star}$, the value update primarily tracks the moving best response $\lambda(\rho_{k})$, so the policy-induced value can stabilize earlier than the primal iterate over the finite training horizon. Panel (b) tracks the initial-state value under the dual-iterate policy: after a rapid increase during early training, the curve approaches the initial state value of the optimal regularized policy. The separation between the optimal unregularized value and the value of the optimal regularized policy shows the suboptimality induced by the entropy regularization. Panel (c) compares PGDA-RL against the generator-based benchmark SPD-dMDP (Chen and Wang, 2016) that builds on the LP approach to (unregularized) MDPs. SPD-dMDP improves more quickly early on, whereas PGDA-RL closes the gap over longer training and achieves comparable final accuracy over the shown horizon. A key difference is the sampling model: SPD-dMDP draws transitions from uniformly sampled state-action pairs, while our method relies on trajectory data generated by the current behavior policy. Moreover, SPD-dMDP reports the averaged iterate as output, which yields a visibly smoother curve than the last-iterate behavior shown for PGDA-RL. Finally, panel (d) reports the KL divergence between the learned dual iterate policy and the optimal regularized policy summed uniformly over non-terminal states, $\mathsf{D}(\pi_{r}^{\star}\parallel\pi_{\rho_{k}})=\sum_{s\in\mathcal{S},a\in\mathcal{A}}\pi_{r}^{\star}(a|s)\log\left(\frac{\pi_{r}^{\star}(a|s)}{\pi_{\rho_{k}}(a|s)}\right)$, which decreases consistently over training, indicating convergence in policy space alongside value improvement. In summary, the simulations support our theoretical findings by demonstrating convergence of the dual-induced policy $\pi_{\rho_{k}}$ toward the unique optimal regularized policy $\pi_{r}^{\star}$ on all non-terminal states.